OpenStudy (anonymous):

-20=-4x-6 how do i solve it?

OpenStudy (anonymous):

-20=-4x-6 -14=-4x -14/-4=x -3.5=x

OpenStudy (anonymous):

you wouldn't subtract the -4x to the -20... combine like terms?

OpenStudy (anonymous):

nope, you have to isolate the x

OpenStudy (anonymous):

meaning get all the numbers to one side of the equation

OpenStudy (anonymous):

and then solve for x, in this example I brought the negative 6 to the otherside. since it switches side, the negative turns into a +6

OpenStudy (anonymous):

giving us -14=-4x then we send the -4 to the otherside by dividing both sides

OpenStudy (anonymous):

Add 6 to both sides of the equation to isolate the X value Then divide both sides by -4 to find X